OpenStudy (anonymous):

David drove 168 miles on 7 gallons of gasoline. Which of the following equations can be used to find how far can he can travel on a full tank of 16 gallons

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

can you help please i don't understand this

OpenStudy (studygurl14):

\(\large\frac{168}{7}=\frac{x}{16}\) solve for x. (cross-multiply)

OpenStudy (anonymous):

well i need to find an equation and the closest one to the one you wrote is 168/7 = 16/x

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

That's incorrect.

OpenStudy (studygurl14):

You can rearrange the formula like this: \(\large\frac{7}{168}=\frac{16}{x}\)

OpenStudy (studygurl14):

You can also write it like this: \(\large\frac{7}{16}=\frac{168}{x}\)